OpenStudy (anonymous):

Jina has scored 22, 24, and 30 points in her three basketball games so far. How many points does she need to score in her next game so that her average (mean) is 26 points per game?

OpenStudy (anonymous):

(22+24+30+x)/n=26

OpenStudy (austinl):

\[\frac{22+24+30+X}{4} = 26\] X is the number of points required to reach that mean.

OpenStudy (anonymous):

Ahhh x=28 (:

OpenStudy (austinl):

Correctomundo